# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 5 Structured and Information-Guided Merging Approaches
While task vector arithmetic and sparsification techniques offer powerful tools for mitigating parameter interference, they primarily operate on static weight representations without considering the underlying model architecture or runtime behavior. We now turn to a complementary class of approaches that leverage structural priors and dynamic information to guide the merging process more effectively. This section examines three advanced paradigms: architecture-level composition methods that preserve modularity through Mixture-of-Experts style routing mechanisms, activation-informed techniques that exploit representation statistics for principled parameter alignment, and evolutionary or search-based meta-strategies that automatically discover optimal merging configurations across the design space.
@@
(Yadav et al., 2024b)
Sparse masking High 2023
@@
(Stoica et al., 2023)
Feature matching N/A Low 2023
@@
(Huang et al., 2023a)
Bayesian composition Partial High 2023
@@
Expert routing N/A High 2025
@@ 5.1 Mixture-of-Experts Style Merging and Expert Routing Strategies
Mixture-of-Experts (MoE) style merging preserves specialized capabilities by organizing distinct expert pathways, directly addressing destructive interference when combining divergent models (Li et al., 2022; Muqeeth et al., 2023). By maintaining separate parameter spaces for each expert and employing learned routing mechanisms, MoE-style merging achieves a principled balance between capability preservation and unified deployment.
The Branch-Train-Merge (BTM) framework provides the conceptual foundation for scalable expert training within this paradigm (Li et al., 2022). In BTM, a shared pretrained LLM is first replicated into multiple branches, each subsequently fine-tuned independently on distinct data subsets or task distributions. The resulting specialized models are then integrated through routing mechanisms that direct inputs to appropriate experts based on learned criteria. Formally, given a set of $K$ expert models $\{E_{1},E_{2},\ldots,E_{K}\}$ derived from a common pretrained initialization, the merged output for an input $x$ is computed as:
$y=\sum_{i=1}^{K}g_{i}(x)\cdot E_{i}(x)$ (8)
@@
with $h(x)$ denoting an intermediate hidden representation and $\{w_{1},\ldots,w_{K}\}$ the learnable routing parameters. To promote computational efficiency (Coste et al., 2024), top-$k$ sparse routing activates only a subset of experts per input:
$g_{i}^{\text{top-}k}(x)=\begin{cases}g_{i}(x)&\text{if }i\in\text{TopK}(\{g_{j}(x)\}_{j=1}^{K})\\ 0&\text{otherwise}\end{cases}$ (10)
@@
enabling inference costs that scale as $O(k/K)$ relative to the full ensemble. CALM (Composition to Augment Language Models) (Bansal et al., 2024) shows that LLMs can be augmented through composition with specialized modules, enabling capability expansion without full model retraining. The WEMoE framework (Tang et al., 2024d) demonstrates how this weight-ensembling approach can effectively integrate multiple task-specific models into a unified MoE architecture, achieving superior multi-task performance compared to naive averaging strategies by constructing experts from merged weight differences rather than full model copies.
@@
Several prominent methods have emerged for constructing MoE systems from independently trained models. PHATGOOSE (Post-Hoc Adaptive Tokenwise Gating Over an Ocean of Specialized Exp.[truncated] addresses the challenge of routing over pre-existing fine-tuned models by learning lightweight gating parameters while keeping expert weights frozen. For each expert $E_{i}$, PHATGOOSE computes routing scores based on the dot product between a learned gate vector $v_{i}$ and the input representation: $g_{i}(x)=v_{i}^{\top}h(x)$, where gate vectors are optimized on each expert’s original training data to identify inputs the expert handles well. MoLE (Mixture of LoRA Experts) (Hu et al., 2022; Feng et al., 2024) takes a parameter-efficient approach by treating individual LoRA adapters as experts, composing the output as $y=W_{0}x+\sum_{i=1}^{K}g_{i}(x)\cdot B_{i}A_{i}x$, where $W_{0}$ represents the frozen pretrained weights and $B_{i}A_{i}$ denotes the low-rank decomposition of expert $i$. This LoRA-as-experts paradigm substantially reduces memory overhead since each expert contributes only the low-rank adapter parameters rather than full model weights. LoRAHub (Huang et al., 2023a) enables efficient cross-task generalization via dynamic LoRA composition, while LoRAMoE (Dou et al., 2023) extends this concept by using LoRA-based experts as plugins that preserve world knowledge during instru.[truncated]
@@
The design of routing mechanisms constitutes a critical component (Muqeeth et al., 2024) determining MoE effectiveness. Gating networks typically employ learned transformations that map input representations to probability distributions over experts. The softmax-based router represents the most common instantiation, computing routing weights as $g(x)=\text{softmax}(W_{g}\cdot h(x))$, where $h(x)$ denotes an intermediate representation and $W_{g}$ represents learnable gating parameters. To promote computational efficiency through sparse activation, top-$k$ routing selects only a subset of experts for each input: $g_{i}(x)=\text{softmax}(\text{TopK}(W_{g}\cdot h(x)))_{i}$, enabling inference costs that scale sublinearly with expert count. Load balancing mechanisms prevent expert collapse, where routing degenerates to selecting only a few experts. The auxiliary load balancing loss $\mathcal{L}_{\text{balance}}=\alpha\cdot K\cdot\sum_{i=1}^{K}f_{i}\cdot p_{i}$ penalizes uneven expert utilization, where $f_{i}$ represents the fraction of tokens routed to expert $i$ and $p_{i}$ denotes the average routing probability for expert $i$ across the batch (Jiang et al., 2024; Fedus et al., 2022). Switch Transformers (Fedus et al., 2022) demonstrated scaling to trillion parameter models with simple and efficient sparsity mechanisms. Branch-Train-MiX (Sukhbaatar et al., 2024) provides a framework for mixing expert LLMs into a unified MoE architecture through branched training procedures. Alternative routing strategies beyond learned softmax gating include expert-choice routing (Jiang et al., 2024), which inverts the selection process by having experts select their top-$k$ preferred tokens rather than tokens selecting experts, naturally achieving load balance. The granularity of routing decisions, whether at token-level for fine-grained adaptation or sequence-level for reduced overhead, depends substantially on expert specialization patterns and input diversity (Yadav et al., 2024a).
Post-hoc integration of specialized fine-tuned models as experts presents unique challenges distinct from training MoE systems from scratch, as experts were never optimized for selective activation. PHATGOOSE addresses this through a two-stage process: first computing per-expert gate vectors by optimizing routing scores on each expert’s training data independently, then combining these gates for joint inference without requiring access to all training data simultaneously. Model Kinship (Hu et al., 2024) provides a framework for exploring and quantifying the relationships between models that influence merging outcomes. Arrow (Yadav et al., 2024a) formulates expert selection as multi-objective optimization, learning to compose subsets of LoRA experts based on task characteristics. ZipIt (Stoica et al., 2023) demonstrates that models trained on different tasks can be merged without additional training through feature matching (Stoica et al., 2023) and permutation alignment. For scenarios where expert training data is unavailable, representation-based heuristics estimate input-expert compatibility through embedding similarity (Hu et al., 2024). Model-Based Experts (Yadav et al., 2024a) further improve efficiency by training small auxiliary classifiers to predict which experts will produce outputs most similar to the full ensemble, enabling dynamic expert pruning at inference time.
@@
MoE-style merging incurs distinct resource tradeoffs compared to parameter-averaging approaches. While parameter averaging produces a single model with memory requirements identical to any constituent, MoE architectures must store all $K$ expert parameter sets, resulting in memory scaling of $O(K\cdot|E|)$ where $|E|$ denotes expert size, though LoRA-based experts substantially mitigate this overhead (Hu et al., 2022). Computationally, sparse top-$k$ routing reduces inference FLOPs to approximately $k/K$ of the dense ensemble cost, but introduces routing computation overhead and memory bandwidth costs from loading selected expert weights. In contrast, merged models via techniques such as TIES-Merging (Yadav et al., 2023) or DARE (Yu et al., 2023) incur no additional inference cost beyond a single model but cannot dynamically adapt computation to input characteristics. The choice between paradigms thus depends on deployment constraints: MoE architectures favor scenarios requiring preserved specialist capabilities and sufficient memory, while parameter averaging suits resource-constrained settings where moderate capability blending suffices. Table 6 compares representative MoE-style and LoRA-based composition methods.
@@
MoLE LoRA adapter $K\times$ LoRA Learned gating 2025 Mixture-of-LoRA-Experts routing
@@
Beyond these structural approaches to combining models, an orthogonal direction leverages empirical observations of how models process inputs to guide the merging process itself.
@@
Despite the architectural elegance, MoE-style merging introduces several practical challenges. The memory footprint scales linearly with the number of experts, as each expert retains a full parameter copy (or full LoRA adapter), making deployment of $K$-expert systems $K$-times more expensive than a single merged model. Router training requires representative held-out data for each task, and poorly calibrated routers may exhibit expert collapse—routing all inputs to a single dominant expert—particularly when task distributions are imbalanced. Li et al. (2022) showed in the Branch-Train-Merge framework that domain classifier accuracy directly bounds the quality of the merged output, establishing a tight coupling between routing quality and downstream performance. Federated variants further exacerbate these issues: heterogeneous client data distributions can cause routers to overfit local patterns (Li et al., 2020).
@@ 5.2 Activation-Informed and Representation-Guided Merging Methods
Activation-informed and representation-guided merging methods embody this empirical approach, representing a significant departure from static weight-averaging by incorporating direct observations of model behavior into the merging process. While methods such as Fisher-weighted averaging utilize second-order gradient information computed at parameter positions, activation-based approaches instead examine how models actually process inputs through their forward computational pathways, thereby capturing functional importance that may not be evident from gradient-based measures alone.
@@
The fundamental insight motivating these approaches is that parameter importance for a given task manifests directly in activation patterns during inference. When a model processes task-relevant inputs, certain neurons and attention heads exhibit consistently high activation magnitudes or frequencies, indicating their centrality to the computational process. Conversely, components with sparse or negligible activations contribute minimally to output formation and thus warrant reduced influence during merging. Formally, given a calibration dataset $\mathcal{D}$ representative of target tasks, activation-informed methods compute importance scores $\alpha_{i}$ for parameter groups by aggregating statistics over forward passes: $\alpha_{i}=f(\{a_{i}(x):x\in\mathcal{D}\})$, where $a_{i}(x)$ denotes the activation vector associated with parameter group $i$ when processing input $x$, and $f$ is an aggregation function such as mean absolute activation or activation variance.
@@
Representation-level alignment extends beyond scalar importance scores to address the geometric structure of intermediate representations across models. Given two models $\theta_{A}$ and $\theta_{B}$ fine-tuned from a common pretrained initialization, their intermediate representations may occupy differently oriented subspaces despite encoding semantically similar information. Centered Kernel Alignment (CKA) and related similarity metrics enable quantification of representational correspondence across layers, providing guidance for matching functionally equivalent components before parameter combination. This alignment proves particularly valuable when models have undergone divergent fine-tuning trajectories that disrupt the implicit correspondence established by shared initialization.
@@
Semantic-guided merging approaches leverage the rich structure of intermediate representations to inform merging decisions at a finer granularity than global importance scores permit. Recent work on multimodal (Lin et al., 2024) large language model (MLLM) merging, as noted in Yang et al. (2024b), demonstrates that semantic information extracted from multimodal representations can guide selective parameter combination in ways that preserve cross-modal alignment. These methods typically construct auxiliary objectives that encourage merged representations to maintain task-relevant semantic properties observed in constituent models, effectively treating representation preservation as a soft constraint during the merging optimization.
Yang et al. (2024b) provide a taxonomy of activation-informed techniques, categorizing them by information source—from lightweight activation magnitudes to sophisticated attention pattern analysis. These methods achieve their advantages through task-relevant calibration data, raising questions about data requirements and calibration set composition.
@@
The computational overhead of activation-informed merging varies substantially across methods. Simple activation magnitude averaging requires only forward passes through calibration data, adding modest preprocessing cost. However, more sophisticated representation alignment techniques may require solving optimization problems over representation matching objectives, substantially increasing computational requirements. Nobari et al. (2025) observe that model merging serves as a transformative method driving substantial advancements when appropriately guided by task-specific information, though the optimal tradeoff between computational investment and merging quality improvement remains context-dependent.
@@
Empirical evidence suggests activation-informed methods demonstrate greatest advantages when merging models with heterogeneous fine-tuning histories or when constituent models exhibit substantial parameter interference under naive averaging. In scenarios where linear mode connectivity (Singh and Jaggi, 2020) holds strongly, typically when models share recent fine-tuning from identical pretrained checkpoints, the marginal benefit of activation guidance diminishes relative to simpler static methods. This aligns with the understanding that shared initialization establishes common loss basins within which parameter averaging succeeds without explicit importance weighting. Consequently, practitioners should consider activation-informed approaches primarily when merging distantly related models or when calibration data clearly reveals differential parameter utilization across tasks. Yet even with principled importance weighting, the broader challenge remains: given a collection of candidate models, how should one select which models to merge, determine layer-wise mixing coefficients, and configure method-specific hyperparameters? This combinatorial complexity motivates automated optimization strategies that can systematically explore the space of possible merge configurations.

### 30 Mar 2026 v2

@@ 5 Structured and Information-Guided Merging Approaches
While task vector arithmetic and sparsification techniques offer powerful tools for mitigating parameter interference, they primarily operate on static weight representations without considering the underlying model architecture or runtime behavior. We now turn to a complementary class of approaches that exploit structural priors and dynamic information to guide the merging process more effectively. This section examines three advanced families of methods. These include architecture-level composition methods that preserve modularity through Mixture-of-Experts (MoE) style routing mechanisms, activation-informed techniques that exploit representation statistics for principled parameter alignment, and evolutionary or search-based meta-strategies that automatically discover optimal merging configurations across the design space.
@@
(Davari and Belilovsky, 2024)
Sparse masking High 2024
@@
(Stoica et al., 2024)
Feature matching N/A Low 2024
@@
(Huang et al., 2023)
Gradient-free composition Partial High 2023
@@
Expert routing N/A High 2024
@@ 5.1 Mixture-of-Experts Style Merging and Expert Routing Strategies
Mixture-of-Experts (MoE) style merging preserves specialized capabilities by organizing distinct expert pathways, directly countering destructive interference when combining divergent models (Li et al., 2022; Muqeeth et al., 2024b). By maintaining separate parameter spaces for each expert and using learned routing mechanisms, MoE-style merging strikes a principled balance between capability preservation and unified deployment.
The Branch-Train-Merge (BTM) framework provides the conceptual foundation for scalable expert training within this approach (Li et al., 2022). In BTM, a shared pretrained LLM is first replicated into multiple branches, each subsequently fine-tuned independently on distinct data subsets or task distributions. The original BTM framework combines expert predictions through simple ensemble averaging at inference time. Subsequent work extended this paradigm by introducing learned routing mechanisms that direct inputs to appropriate experts, yielding a general MoE-style formulation. Formally, given a set of $K$ expert models $\{E_{1},E_{2},\ldots,E_{K}\}$ derived from a common pretrained initialization, the merged output for an input $x$ is computed as:
$y=\sum_{i=1}^{K}g_{i}(x)\cdot E_{i}(x),$ (8)
@@
with $h(x)$ denoting an intermediate hidden representation and $\{w_{1},\ldots,w_{K}\}$ the learnable routing parameters. To promote computational efficiency, top-$k$ sparse routing activates only a subset of experts per input:
$g_{i}^{\text{top-}k}(x)=\begin{cases}g_{i}(x)&\text{if }i\in\text{TopK}(\{g_{j}(x)\}_{j=1}^{K})\\ 0&\text{otherwise}\end{cases},$ (10)
@@
enabling inference costs that scale as $O(k/K)$ relative to the full ensemble. CALM (Composition to Augment Language Models) (Bansal et al., 2024) shows that LLMs can be augmented through composition with specialized modules, allowing capability expansion without full model retraining. The WEMoE framework (Tang et al., 2024e) shows how this weight-ensembling approach can effectively integrate multiple task-specific models into a unified MoE architecture, achieving superior multi-task performance compared to naive averaging strategies by constructing experts from merged weight differences rather than full model copies.
@@
Several prominent methods have been proposed for constructing MoE systems from independently trained models. PHATGOOSE (Post-Hoc Adaptive Tokenwise Gating Over an Ocean of Specialized Experts) (Muqeeth et al., 2024a) addresses the challenge of routing over pre-existing fine-tuned models by learning lightweight gating parameters while keeping expert weights frozen. For each expert $E_{i}$, PHATGOOSE computes routing scores based on the dot product between a learned gate vector $v_{i}$ and the input representation: $g_{i}(x)=v_{i}^{\top}h(x)$, where gate vectors are optimized on each expert’s original training data to identify inputs the expert handles well. MoLE (Mixture of LoRA Experts) (Wu et al., 2024) takes a parameter-efficient approach by treating individual LoRA adapters as experts, composing the output as $y=W_{0}x+\sum_{i=1}^{K}g_{i}(x)\cdot B_{i}A_{i}x$, where $W_{0}$ represents the frozen pretrained weights and $B_{i}A_{i}$ denotes the low-rank decomposition of expert ii. This LoRA-as-experts design greatly reduces memory overhead since each expert contributes only the low-rank adapter parameters rather than full model weights. LoRAHub (Huang et al., 2023) enables efficient cross-task generalization via dynamic LoRA composition, while LoRAMoE (Dou et al., 2024) extends this concept by using LoRA-based experts as plugins that preserve world knowledge during instruction tuning.
@@
Beyond the basic softmax gating (Eq. 9) and top-$k$ routing (Eq. 10) introduced above, several design choices critically determine MoE effectiveness. Load balancing mechanisms prevent expert collapse, where routing degenerates to selecting only a few experts. The auxiliary load balancing loss $\mathcal{L}_{\text{balance}}=\alpha\cdot K\cdot\sum_{i=1}^{K}f_{i}\cdot p_{i}$ penalizes uneven expert usage, where $f_{i}$ represents the fraction of tokens routed to expert $i$ and $p_{i}$ denotes the average routing probability for expert $i$ across the batch (Jiang et al., 2024; Fedus et al., 2022). Switch Transformers (Fedus et al., 2022) scaled to trillion parameter models with simple and efficient sparsity mechanisms. Branch-Train-MiX (Sukhbaatar et al., 2024) provides a framework for mixing expert LLMs into a unified MoE architecture through branched training procedures. Alternative routing strategies beyond learned softmax gating include expert-choice routing, which inverts the selection process by having experts select their top-$k$ preferred tokens rather than tokens selecting experts, naturally achieving load balance. The granularity of routing decisions, whether at token-level for fine-grained adaptation or sequence-level for reduced overhead, depends considerably on expert specialization patterns and input diversity (Yadav et al., 2025a).
Post-hoc integration of specialized fine-tuned models as experts presents unique challenges distinct from training MoE systems from scratch, as experts were never optimized for selective activation. PHATGOOSE tackles this through a two-stage process, first computing per-expert gate vectors by optimizing routing scores on each expert’s training data independently, then combining these gates for joint inference without requiring access to all training data simultaneously. Merge, Then Compress (Li et al., 2024b) demystifies efficient sparse MoE by first merging redundant experts and then compressing the resulting architecture, providing insights into the relationship between expert similarity and routing behavior. LoRA-Flow (Wang et al., 2024) introduces dynamic LoRA fusion that adaptively combines LoRA adapters during generation based on input characteristics, allowing more fine-grained expert use than static routing. LoraRetriever (Zhao et al., 2024b) proposes input-aware LoRA retrieval and composition for mixed tasks, dynamically selecting and combining relevant adapters from a library. XFT (Ding et al., 2024) unlocks the power of code instruction tuning by merging upcycled mixture-of-experts, showing that MoE structures can emerge from strategic combination of specialized models. Model Kinship (Hu et al., 2025b) provides a framework for exploring and quantifying the similarity relationships between LLMs that predict merging performance, supporting principled selection of candidate models for merging through a kinship-based greedy strategy.
@@
MoE-style merging incurs distinct resource tradeoffs compared to parameter-averaging approaches. While parameter averaging produces a single model with memory requirements identical to any constituent, MoE architectures must store all $K$ expert parameter sets, resulting in memory scaling of $O(K\cdot|E|)$ where $|E|$ denotes expert size. LoRA-based experts greatly mitigate this overhead (Hu et al., 2022), and SMILE (Tang et al., 2024b) further reduces costs by constructing sparse mixture of low-rank experts from pretrained models in a zero-shot manner. Computationally, sparse top-$k$ routing reduces inference FLOPs to approximately $k/K$ of the dense ensemble cost but introduces routing overhead and memory bandwidth costs from loading selected expert weights. In contrast, merged models via TIES-Merging (Yadav et al., 2023) or DARE (Yu et al., 2024) incur no additional inference cost beyond a single model but cannot dynamically adapt computation to input characteristics.
Several recent methods further improve LoRA-based MoE composition. LoRA Soups (Prabhakar et al., 2025) show that model merging outperforms data mixing for skill composition, with their concatenation-of-LoRAs (CAT) method achieving optimal weighting of individually trained modules. MergeME (Zhou et al., 2025) extends merging techniques to both homogeneous and heterogeneous MoE architectures for growing MoE-based LLM deployments. DLP-LoRA (Zhang and Li, 2024) introduces dynamic lightweight plugins for efficient task-specific fusion, while Decouple and Orthogonalize (Zheng et al., 2025b) proposes data-free LoRA merging through orthogonal subspace decomposition, and Zhang and Zhou (2025) analyze LoRA interference patterns and propose orthogonal subspace methods for robust merging. The choice between MoE and parameter-averaging approaches thus depends on deployment constraints. Specifically, MoE architectures favor scenarios requiring preserved specialist capabilities and sufficient memory, while parameter averaging suits resource-constrained settings where moderate capability blending suffices. Table 6 compares representative MoE-style and LoRA-based composition methods.
@@
MoLE
(Wu et al., 2024)
LoRA adapter $K\times$ LoRA Learned gating 2024 Mixture-of-LoRA-Experts routing
@@
Beyond these structural approaches to combining models, an orthogonal direction exploits empirical observations of how models process inputs to guide the merging process itself.
@@
Despite the architectural elegance, MoE-style merging introduces several practical challenges. The memory footprint scales linearly with the number of experts, as each expert retains a full parameter copy (or full LoRA adapter), making deployment of $K$-expert systems $K$-times more expensive than a single merged model. Router training requires representative held-out data for each task, and poorly calibrated routers may exhibit expert collapse, where the router directs all inputs to a single dominant expert, particularly when task distributions are imbalanced. Li et al. (2022) showed in the Branch-Train-Merge framework that expert domain specialization is essential for effective merging, as random data splits fail to produce well-performing merged models. Federated variants further exacerbate these issues, since heterogeneous client data distributions can cause routers to overfit local patterns (Li et al., 2020).
@@ 5.2 Activation-Informed and Representation-Guided Merging Methods
Activation-informed and representation-guided merging methods embody this empirical approach, representing a departure from static weight-averaging by incorporating direct observations of model behavior into the merging process. While methods such as Fisher-weighted averaging use second-order gradient information computed at parameter positions, activation-based approaches instead examine how models actually process inputs through their forward computational pathways, thereby capturing functional importance that may not be evident from gradient-based measures alone.
@@
A key insight motivating these approaches is that parameter importance for a given task manifests directly in activation patterns during inference. When a model processes task-relevant inputs, certain neurons and attention heads exhibit consistently high activation magnitudes or frequencies, indicating their centrality to the computational process. Conversely, components with sparse or negligible activations contribute minimally to output formation and thus warrant reduced influence during merging. Formally, given a calibration dataset $\mathcal{D}$ representative of target tasks, activation-informed methods compute importance scores $\alpha_{i}$ for parameter groups by aggregating statistics over forward passes: $\alpha_{i}=f(\{a_{i}(x):x\in\mathcal{D}\})$, where $a_{i}(x)$ denotes the activation vector associated with parameter group $i$ when processing input $x$, and $f$ is an aggregation function such as mean absolute activation or activation variance.
@@
Representation-level alignment extends beyond scalar importance scores to account for the geometric structure of intermediate representations across models. Given two models $\theta_{A}$ and $\theta_{B}$ fine-tuned from a common pretrained initialization, their intermediate representations may occupy differently oriented subspaces despite encoding semantically similar information. Centered Kernel Alignment (CKA) and related similarity metrics can quantify representational correspondence across layers, providing guidance for matching functionally equivalent components before parameter combination. This alignment proves particularly valuable when models have undergone divergent fine-tuning trajectories that disrupt the implicit correspondence established by shared initialization.
@@
Semantic-guided merging approaches exploit the rich structure of intermediate representations to inform merging decisions at a finer granularity than global importance scores permit. Multimodal model development has increasingly adopted MoE-style architectures, with MoE-LLaVA (Lin et al., 2024) showing that sparse MoE training strategies can effectively scale large vision-language models while maintaining computational efficiency. More broadly, merging techniques applied to multimodal models must account for cross-modal alignment, as parameter combination may disrupt the correspondence between vision and language representations established during joint training. These methods typically construct auxiliary objectives that encourage merged representations to maintain task-relevant semantic properties observed in constituent models, effectively treating representation preservation as a soft constraint during the merging optimization.
Yang et al. (2024b) provide a taxonomy of activation-informed techniques, categorizing them by information source, ranging from lightweight activation magnitudes to detailed attention pattern analysis. Activated Parameter Locating (Kong et al., 2024) introduces causal intervention to identify task-relevant parameters, providing a principled mechanism for determining which parameters should be prioritized during merging. Activation-Guided Consensus Merging (Yao et al., 2025) extends this direction by computing consensus across multiple models’ activation patterns to guide parameter combination decisions. Curvature-informed merging (Mahdavinia et al., 2025) harnesses optimization dynamics to incorporate second-order curvature information into the merging process, providing a bridge between activation-based and gradient-based importance estimation. These methods achieve their advantages through task-relevant calibration data, raising questions about data requirements and calibration set composition.
@@
Computational overhead of activation-informed merging varies substantially across methods. Simple activation magnitude averaging requires only forward passes through calibration data, adding modest preprocessing cost. However, more involved representation alignment techniques may require solving optimization problems over representation matching objectives, considerably increasing computational requirements. The optimal tradeoff between computational investment and merging quality improvement remains context-dependent (Nobari et al., 2025).
@@
Empirical evidence suggests activation-informed methods show the greatest advantages when merging models with heterogeneous fine-tuning histories or when constituent models exhibit substantial parameter interference under naive averaging. In scenarios where linear mode connectivity (Frankle et al., 2020) holds strongly, typically when models share recent fine-tuning from identical pretrained checkpoints, the marginal benefit of activation guidance diminishes relative to simpler static methods. This aligns with the understanding that shared initialization establishes common loss basins within which parameter averaging succeeds without explicit importance weighting. Consequently, practitioners should consider activation-informed approaches primarily when merging distantly related models or when calibration data clearly reveals differential parameter usage across tasks. Yet even with principled importance weighting, the broader challenge remains. Given a collection of candidate models, how should one select which models to merge, determine layer-wise mixing coefficients, and configure method-specific hyperparameters? This combinatorial complexity motivates automated optimization strategies that can systematically explore the space of possible merge configurations.